OpenStudy (anonymous):

how many solutions does y=1/2x+2 and 4x-8y=-8

OpenStudy (johnweldon1993):

to do this...let us set them equal to each other y = 1/2x + 2 4x - 8y = -8 hmm..that second equation look s abit messy..lets turn in into slope intercept form...y=mx + b -8y = -8 - 4x y = 1 + 4/8x or y = 1 + 1/2x written correctly to match with the equation y = 1/2x + 1 so we have 2 equations now y = 1/2x + 1 and y = 1/2x + 2 hmm same slope huh? odd isn't it...if a line has the SAME slope but different y-intercept as another line....what does that look like? do they ever meet?